OpenStudy (anonymous):

determine the maximum and minimum values for the equation : y=2x^3/3+X^2-6x+x

OpenStudy (asnaseer):

the general procedure you should follow is: 1) you need to find the first differential of y with respect to x, 2) then find what values of x give a zero value for the first differential, 3) then plug each of these x values into the second differential 3.1) if the 2nd differential is negative, then you have a local maximum, 3.2) if its positive, then you have a local minimum, 3.3) if its zero, then you have a point of inflection. 4) plug each x value into the original equation to find the y value at the maximum, minimum, or point of inflection so, with your equation we get:\[\begin{align} y&=\frac{2x^3}{3}+x^2-6x+1=\frac{2x^3}{3}+x^2-5x\\ \therefore y'&=\frac{2*3x^2}{3}+2x-6=2x^2+2x-5\\ &\text{following step 2 above, we need to find which values}\\ &\text{of x make y' zero, so:}\\ 0&=2x^2+2x-5\\ &\text{solving this the quadratic equation gives:}\\ x&=\frac{-1\pm\sqrt{11}}{2}\\ &\text{we now need to work out the 2nd derivative for step 3 above:}\\ y''&=4x+2\\ &\text{this gives a positive value for } x=\frac{-1+\sqrt{11}}{2}\implies min\\ &\text{and gives a negative value for }x=\frac{-1-\sqrt{11}}{2}\implies max\\ \end{align}\]so now you can follow step 4 above to find the minimum and maximum values of y at each of these x values.

OpenStudy (lollylau):

Do we actually NEED to use 2nd derivatives??

OpenStudy (asnaseer):

The 1st derivative is used to find the points at which the slope is zero. The 2nd derivative tells you what each of these points represent - i.e. are they maximum, minimum or points of inflection.